OpenStudy (anonymous):

What is the given term in each geometric sequence? 1.) 6th term: 9, 18, 36,... 2.) 5th term: -3, -15, -75,...

OpenStudy (maddyx):

The common difference is 2. So, nth term of Geometric sequence is = a*r^(n-1) 1) 6th term = 9*(2^(6-1)) = 288 2) 5th term = (-3)*(5^(5-1)) = -1875